solve for x
0.25mg: IML :: 625 mcg : XML 0 100 100 10 al al 000 roll mboo N -
Answer #1

ANSWER.

* 0.25 mg = 1 ml   , 62.5 mcg = X ml.

* Convert 0.25 mg to mcg for calculation.

             1 mg = 1000 mcg.

             0.25 mg = 250 mcg.

* 250 mcg = 1 ml. , 62.5 mcg = X ml.

           250 mcg = 1 ml

           62.5 mcg = X ml

        X * 250 = 1 * 62.5

             X = 1 * 62.5 / 250

                 = 0.25 ml.